# Generalized $T\bar{T}$-like flows for scalar theories in two dimensions

### v1

@@ 3 General irrelevant $T\bar{T}$-like deformations for perturbative solutions
We investigate the perturbative solutions of the PDE 1.10, described by general irrelevant $T\bar{T}$-like deformations. This type has a dimensionful $\lambda$ coupling, and we take it as a function of the integers power of $P_{1}$ and $P_{2}$ in the form of $\mathcal{L}(\lambda^{n})=\mathcal{K}(P_{1}^{N},P_{2}^{M},\lambda^{N+2M-1})$. The Lagrangian can be expressed up to the $\lambda^{7}$ order, incorporating arbitrary coefficients denoted by $a_{n}$, in the following manner:
@@
$+\lambda^{4}\bigl{(}a_{9}{}{P_{1}}^{5}+a_{10}{}{P_{1}}^{3}P_{2}{}+a_{11}{}P_{1}{}{P_{2}}^{2}\bigr{)}+\lambda^{5}\bigl{(}a_{12}{}{P_{1}}^{6}+a_{13}{}{P_{1}}^{4}P_{2}{}+a_{14}{}{P_{1}}^{2}{P_{2}}^{2}+a_{15}{}{P_{2}}^{3}\bigr{)}$
@@
This section aims to solve the PDE sequentially 1.10, utilizing the Lagrangian of order $\lambda^{7}$. This process will clarify the relationship between the coefficients $a_{n}$ at each order of $\lambda$, ultimately leading to the following correlation among the coefficients $a_{n}$. This correlation is detailed in Appendix A. Upon applying conditions A to the Lagrangian 3.1, a solution to equation 1.10 emerges, involving several constants $a_{2}$, $a_{7}$, $a_{13}$ and $a_{21}$ which remain undetermined. At the order of $\lambda^{n}$, $\frac{n}{2}$ coefficients are not fixed. This general Lagrangian is obtained as follows:
${\cal L}(\lambda)$ $=$ $-\tfrac{1}{2}P_{1}{}+\lambda a_{2}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}+4\lambda^{2}{a_{2}}^{2}P_{1}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}+\lambda^{3}\Bigl{(}-\tfrac{1}{4}a_{7}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}+8{a_{2}}^{3}\bigl{(}{P_{1}}^{4}-4{P_{2}}^{2}\bigr{)}\Bigr{)}$ (3.2)
@@
$+\lambda^{5}\Bigl{(}-\tfrac{1}{6}a_{13}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}-\tfrac{40}{3}{a_{2}}^{2}a_{7}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}{P_{1}}^{2}+4P_{2}{}\bigr{)}-\tfrac{512}{3}{a_{2}}^{5}\bigl{(}{P_{1}}^{6}-18{P_{1}}^{2}{P_{2}}^{2}+28{P_{2}}^{3}\bigr{)}\Bigr{)}$
@@
$+\lambda^{7}\Bigl{(}-\tfrac{1}{8}a_{21}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{4}+7a_{2}{}{a_{7}}^{2}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}\bigl{(}{P_{1}}^{2}+6P_{2}{}\bigr{)}-14{a_{2}}^{2}a_{13}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}\bigl{(}{P_{1}}^{2}+6P_{2}{}\bigr{)}$
$+448{a_{2}}^{4}a_{7}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}3{P_{1}}^{4}+12{P_{1}}^{2}P_{2}{}-56{P_{2}}^{2}\bigr{)}+4096{a_{2}}^{7}\bigl{(}7{P_{1}}^{8}-206{P_{1}}^{4}{P_{2}}^{2}+608{P_{1}}^{2}{P_{2}}^{3}-504{P_{2}}^{4}\bigr{)}\Bigr{)}$
@@
We aim to derive the flow equations pertinent to this overarching Lagrangian in 3.2. Such equations will conform to the archetype of standard irrelevant $T\bar{T}$ deformations within a two-dimensional framework, applicable to theories characterized by integer powers of the Lorentz Invariant variables $P_{1}$ and $P_{2}$. To accomplish this, we commence by determining the energy-momentum tensor from the general Lagrangian, denoted as Eq. 3.2, derived in the following manner:
@@
Appendix A contains the details of the two functions ${\cal F}(P_{1},P_{2},\lambda)$ and ${\cal G}(P_{1},P_{2},\lambda)$, specifically in Eqs. A and A. In the context of the various theories governed by PDE 1.10, the flow equation is contingent upon two specific variables, denoted as $T_{\mu\nu}T^{\mu\nu}$ and $T^{\mu}{}_{\mu}T^{\nu}{}_{\nu}$. To elucidate, the flow equation is generally represented by the formula $\frac{\partial{{\cal L}(\lambda)_{NGZ}}}{\partial{\lambda}}=f(T_{\mu\nu}T^{\mu\nu},T^{\mu}{}_{\mu}T^{\nu}{}_{\nu})$. For the derivation of $f(T_{\mu\nu}T^{\mu\nu},T^{\mu}{}_{\mu}T^{\nu}{}_{\nu})$, it is imperative to construct two foundational structures: $T_{\mu\nu}T^{\mu\nu}$ and $T^{\mu}{}_{\mu}T^{\nu}{}_{\nu}$. The configuration of $T_{\mu\nu}T^{\mu\nu}$ can be deduced from 3.3 via the equation:
@@
$+\lambda^{4}\Bigl{(}a_{2}{}a_{7}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}29{P_{1}}^{2}+22P_{2}{}\bigr{)}+32{a_{2}}^{4}\bigl{(}39{P_{1}}^{6}-174{P_{1}}^{4}P_{2}{}+148{P_{1}}^{2}{P_{2}}^{2}+88{P_{2}}^{3}\bigr{)}\Bigr{)}$
@@
$+\lambda^{7}\Bigl{(}32a_{2}{}{a_{7}}^{2}P_{1}{}\bigl{(}3{P_{1}}^{2}-34P_{2}{}\bigr{)}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}-\tfrac{256}{3}{a_{2}}^{2}a_{13}{}P_{1}{}\bigl{(}2{P_{1}}^{2}-25P_{2}{}\bigr{)}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}$
@@
have been employed. The structure of $T^{\mu}{}_{\mu}T^{\nu}{}_{\nu}$ can be directly derived from 3.3, as delineated below:
@@
$+\lambda^{6}\Bigl{(}\tfrac{9}{4}{a_{7}}^{2}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{4}-\tfrac{20}{3}a_{2}{}a_{13}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{4}-\tfrac{16}{3}{a_{2}}^{3}a_{7}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}\bigl{(}319{P_{1}}^{2}+454P_{2}{}\bigr{)}$
@@
$-\tfrac{736}{3}{a_{2}}^{2}a_{13}{}P_{1}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{4}+\tfrac{2048}{3}{a_{2}}^{7}P_{1}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}415{P_{1}}^{4}-3004{P_{1}}^{2}P_{2}{}+5692{P_{2}}^{2}\bigr{)}\Bigr{)}$
In our previous work, detailed in [50], we established that within a four-dimensional framework, the general form of irrelevant $T\bar{T}$-like deformations for actions characterized by integral exponents of the Lorentzian variables $t$ and $z$ manifests as a sequence comprising dual constructs: $T_{\mu\nu}T^{\mu\nu}$ and $T_{\mu}{}^{\mu}{}T_{\nu}{}^{\nu}$. This pattern is also discernible within a two-dimensional context for the overarching Lagrangian labeled as 3.2. Consequently, we can deduce the expression for this universal distortion, articulated as the ensuing general flow equation, accurate to the order of $\lambda^{7}$:
@@
$c_{0}=-2a_{2},\,\,\,\,\,c_{1}=-\frac{16{a_{2}}^{3}+3a_{7}{}}{16{a_{2}}^{2}},\,\,\,\,\,c_{2}=-\frac{50176{a_{2}}^{6}+1664{a_{2}}^{3}a_{7}{}-27{a_{7}}^{2}-20a_{2}{}a_{13}{}}{768{a_{2}}^{5}},$ (3.8)
$c_{3}=-\frac{121044992{a_{2}}^{9}+4264960{a_{2}}^{6}a_{7}{}-40512{a_{2}}^{3}{a_{7}}^{2}+567{a_{7}}^{3}-38656{a_{2}}^{4}a_{13}{}+720a_{2}{}a_{7}{}a_{13}{}+168{a_{2}}^{2}a_{21}{}}{49152{a_{2}}^{8}}.$
Flow equation 3.7 represents the universal formulation of flow dynamics in $D=2$, encompassing all theoretical models with the appropriate exponents of $P_{1}$ and $P_{2}$, as they pertain to PDE 1.10. This particular flow equation aligns with the overarching principles of the general Lagrangian, denoted as equation 3.2.
@@ 4 General root $T\bar{T}$-like flows in two dimensions
The duality-invariant PDE 1.10 can be simplified by expressing $\mathcal{L}$ as a function of $P_{1}$ and
@@
Upon implementing the variable substitution 4.1, the GSMM Lagrangian in 1.7 is reformulated thus:
@@
It can be explicitly verified that Lagrangian 4.3 satisfies PDE 4.2. A similar approach to the previous section can solve self-duality PDE 4.2 with a series of integer powers of the two independent variables, $P_{1}$ and $\mathcal{P}$. In all the theories we consider, there are only two independent invariants $P_{1}$ and $\mathcal{P}$. All higher-order invariants are combinations of these two. For this purpose, we consider the following generic Lagrangian:
@@
By employing a perturbation approach similar to the previous section, we can solve PDE 4.2 using the general solution 4.4, step by step. These conditions are essential for solving Eq. 4.2 up to the $\lambda$ order in the form of $\Big{(}d_{2}{}\to-\tfrac{1}{2}\sqrt{1+4{d_{1}}^{2}}\Big{)}\,\,\,\,\,,\,\,\,\,\,\,\Big{(}d_{3}\to-\frac{d_{4}{}}{4{d_{1}}}\sqrt{1+4{d_{1}}^{2}}\,\,,\,\,\,d_{5}\to-\frac{d_{1}d_{4}}{\sqrt{1+4{d_{1}}^{2}}}\Big{)}$. In this scenario, only one unknown coefficient $d_{n}$ remains at each order of $\lambda$. The perturbed solutions of PDE 4.2, assuming $d_{1}=-\tfrac{1}{2}\sinh(\gamma)$, up to order of $\lambda^{4}$ are as follows:
$\mathcal{L}(\lambda,\gamma)$ $=$ $-\tfrac{1}{2}\bigl{(}P_{1}\cosh(\gamma)+\mathcal{P}\sinh(\gamma)\bigr{)}+\lambda\,d_{4}\frac{\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{2}}{2\coth(\gamma)}$ (4.5)
$+\lambda^{2}\Bigl{(}\tfrac{1}{3}d_{4}^{2}\bigl{(}\mathcal{P}\mathrm{csch}(\gamma)-2P_{1}\mathrm{sech}(\gamma)\bigr{)}\bigl{(}\mathcal{P}\mathrm{csch}(\gamma)+P_{1}\mathrm{sech}(\gamma)\bigr{)}^{2}+\frac{d_{7}\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{3}}{3\bigl{(}\coth(\gamma)\bigr{)}^{2}}\Bigr{)}$
$+\lambda^{3}\biggl{(}\frac{d_{4}d_{7}\bigl{(}\mathcal{P}\mathrm{csch}(\gamma)-3P_{1}\mathrm{sech}(\gamma)\bigr{)}\bigl{(}\mathcal{P}\mathrm{csch}(\gamma)+P_{1}\mathrm{sech}(\gamma)\bigr{)}^{3}}{2\mathrm{csch}(\gamma)}+\frac{d_{11}\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{4}}{4\bigl{(}\coth(\gamma)\bigr{)}^{3}}$
@@
×$\times\Bigl{(}-68P_{1}^{2}+9\mathcal{P}^{2}+4(20P_{1}^{2}+3\mathcal{P}^{2})\cosh(2\gamma)+3(-4P_{1}^{2}+\mathcal{P}^{2})\cosh(4\gamma)-72P_{1}\mathcal{P}\bigl{(}\cosh(\gamma)\bigr{)}^{3}\sinh(\gamma)\Bigr{)}$
@@
The energy-momentum tensor corresponding to the general Lagrangian 4.5 can be determined as follows:
@@
The functions ${\cal H}(P_{1},\mathcal{P},\lambda,\,\gamma)$ and ${\cal U}(P_{1},\mathcal{P},\lambda,\,\gamma)$ are defined in Appendix A , in Eqs. A.4 and A. To derive a root flow equation, we construct two configurations, $T_{\mu\nu}T^{\mu\nu}$ and $T^{\mu}{}_{\mu}T^{\nu}{}_{\nu}$, from the energy-momentum tensor presented in Eq. 4.6. The configurations for these two structures are provided in Appendix A, in Eqs. A.6 and A.7. We aim to study the root flow equation within the framework of the general Lagrangian 4.5. We consider a general form of the root flow equation derived from the irrelevant flow equation 3.7. Consequently, the general Lagrangian 4.5 must satisfy the following flow equation:
@@
The general Lagrangian 4.5 simplifies to the (SMM) theory in 1.6 in the limit of $\lambda=0$. Notably, there is no $\gamma$-dependent coefficient. The solution to root flow equation 4.7 is derived with the constant $e_{0}=\tfrac{1}{2}$. In the next order of $\lambda$, we encounter the unknown constant $d_{4}$, which determines the $\gamma$-dependence of $d_{4}$ from the solution of root flow equation 4.7. To achieve this, we substitute A.6, A.7, and 4.5 into root flow equation 4.7 and simplify in the order of $\lambda$, resulting in the following differential equation:
$0\displaystyle 2\cosh(2\gamma){d}_{4}-\sinh(2\gamma){d}^{\prime}_{4}=$ 0 . (4.8)
Solving differential equation 4.8 yields the $\gamma$-dependence of the coefficient $d_{4}$ as follows:
@@
where $n_{1}$ is a constant independent of $\gamma$. By applying this method to the order of $\lambda^{2}$, solving root flow equation 4.7 results in a differential equation that depends on the coefficient $d_{7}$ and ${d}^{\prime}_{7}$. The differential equation is as follows:
$n_{1}^{2}\bigl{(}16\cosh(2\gamma)+3(1+4e_{1})\cosh(4\gamma)-51-12e_{1}\bigr{)}+d_{7}\bigl{(}64\mathrm{sech}(\gamma)-96\cosh(\gamma)\bigr{)}+32{d}^{\prime}_{7}\sinh(\gamma)=0.$ (4.10)
By solving differential equation 4.10, we obtain the $\gamma$-dependence of the coefficient $d_{7}$ as follows:
@@
${d}_{11}$ $=$ $\tfrac{1}{2}n_{1}^{3}\cosh(\gamma)\bigl{(}3(1+4e_{1})\gamma\cosh(3\gamma)-8\sinh(\gamma)\bigr{)}+\bigl{(}n_{3}+n_{1}^{3}(\gamma+4e_{1}\gamma)^{2}\bigr{)}\bigl{(}\cosh(\gamma)\bigr{)}^{3}\sinh(\gamma)$ (4.12)

### v2

@@ 3 General irrelevant $T\bar{T}$-like deformations for perturbative solutions
We investigate the perturbative solutions of the PDE (1.10), described by general irrelevant $T\bar{T}$-like deformations. This type has a dimensionful $\lambda$ coupling, and we take it as a function of the integers power of $P_{1}$ and $P_{2}$ in the form of $\mathcal{L}(\lambda^{n})=\mathcal{K}(P_{1}^{N},P_{2}^{M},\lambda^{N+2M-1})$. The Lagrangian can be expressed up to the $\lambda^{7}$ order, incorporating arbitrary coefficients denoted by $a_{n}$, in the following manner:
@@
$+\lambda^{4}\bigl{(}a_{9}{}{P_{1}}^{5}+a_{10}{}{P_{1}}^{3}P_{2}{}+a_{11}{}P_{1}{}{P_{2}}^{2}\bigr{)}$
$+\lambda^{5}\bigl{(}a_{12}{}{P_{1}}^{6}+a_{13}{}{P_{1}}^{4}P_{2}{}+a_{14}{}{P_{1}}^{2}{P_{2}}^{2}+a_{15}{}{P_{2}}^{3}\bigr{)}$
@@
This section aims to solve the PDE sequentially (1.10), utilizing the Lagrangian of order $\lambda^{7}$. This process will clarify the relationship between the coefficients $a_{n}$ at each order of $\lambda$, ultimately leading to the following correlation among the coefficients $a_{n}$. This correlation is detailed in Appendix (A). Upon applying conditions (A.1) to the Lagrangian (3.1), a solution to equation (1.10) emerges, involving several constants $a_{2}$, $a_{7}$, $a_{13}$ and $a_{21}$ which remain undetermined. At the order of $\lambda^{n}$, $\frac{n}{2}$ coefficients are not fixed. This general Lagrangian is obtained as follows:
${\cal L}(\lambda)$ $=$ $-\tfrac{1}{2}P_{1}{}+\lambda a_{2}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}+4\lambda^{2}{a_{2}}^{2}P_{1}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}$
$+\lambda^{3}\Bigl{(}-\tfrac{1}{4}a_{7}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}+8{a_{2}}^{3}\bigl{(}{P_{1}}^{4}-4{P_{2}}^{2}\bigr{)}\Bigr{)}$
@@
$+\lambda^{5}\Bigl{(}-\tfrac{1}{6}a_{13}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}-\tfrac{40}{3}{a_{2}}^{2}a_{7}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}{P_{1}}^{2}+4P_{2}{}\bigr{)}$
$-\tfrac{512}{3}{a_{2}}^{5}\bigl{(}{P_{1}}^{6}-18{P_{1}}^{2}{P_{2}}^{2}+28{P_{2}}^{3}\bigr{)}\Bigr{)}$
@@
$+\lambda^{7}\Bigl{(}-\tfrac{1}{8}a_{21}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{4}+7a_{2}{}{a_{7}}^{2}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}\bigl{(}{P_{1}}^{2}+6P_{2}{}\bigr{)}$
$-14{a_{2}}^{2}a_{13}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}\bigl{(}{P_{1}}^{2}+6P_{2}{}\bigr{)}+448{a_{2}}^{4}a_{7}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}3{P_{1}}^{4}+12{P_{1}}^{2}P_{2}{}-56{P_{2}}^{2}\bigr{)}$
$+4096{a_{2}}^{7}\bigl{(}7{P_{1}}^{8}-206{P_{1}}^{4}{P_{2}}^{2}+608{P_{1}}^{2}{P_{2}}^{3}-504{P_{2}}^{4}\bigr{)}\Bigr{)}$
@@
We aim to derive the flow equations pertinent to this overarching Lagrangian in (3). Such equations will conform to the archetype of standard irrelevant $T\bar{T}$ deformations within a two-dimensional framework, applicable to theories characterized by integer powers of the Lorentz Invariant variables $P_{1}$ and $P_{2}$. To accomplish this, we commence by determining the energy-momentum tensor from the general Lagrangian, denoted as Eq. (3), derived in the following manner:
@@
[…]
@@
$+\lambda^{4}\Bigl{(}a_{2}{}a_{7}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}29{P_{1}}^{2}+22P_{2}{}\bigr{)}$
$+32{a_{2}}^{4}\bigl{(}39{P_{1}}^{6}-174{P_{1}}^{4}P_{2}{}+148{P_{1}}^{2}{P_{2}}^{2}+88{P_{2}}^{3}\bigr{)}\Bigr{)}$
@@
$+\lambda^{7}\Bigl{(}32a_{2}{}{a_{7}}^{2}P_{1}{}\bigl{(}3{P_{1}}^{2}-34P_{2}{}\bigr{)}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}$
$-\tfrac{256}{3}{a_{2}}^{2}a_{13}{}P_{1}{}\bigl{(}2{P_{1}}^{2}-25P_{2}{}\bigr{)}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}$
@@
have been employed. The structure of $T^{\mu}{}_{\mu}T^{\nu}{}_{\nu}$ can be directly derived from (3.3), as delineated below:
@@
$+\lambda^{6}\Bigl{(}\tfrac{9}{4}{a_{7}}^{2}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{4}-\tfrac{20}{3}a_{2}{}a_{13}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{4}$
$-\tfrac{16}{3}{a_{2}}^{3}a_{7}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{3}\bigl{(}319{P_{1}}^{2}+454P_{2}{}\bigr{)}$
@@
$-\tfrac{736}{3}{a_{2}}^{2}a_{13}{}P_{1}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{4}$
$+\tfrac{2048}{3}{a_{2}}^{7}P_{1}{}\bigl{(}{P_{1}}^{2}-2P_{2}{}\bigr{)}^{2}\bigl{(}415{P_{1}}^{4}-3004{P_{1}}^{2}P_{2}{}+5692{P_{2}}^{2}\bigr{)}\Bigr{)}.$
In our previous work, detailed in [55], we established that within a four-dimensional framework, the general form of irrelevant $T\bar{T}$-like deformations for actions characterized by integral exponents of the Lorentzian variables $t$ and $z$ manifests as a sequence comprising dual constructs: $T_{\mu\nu}T^{\mu\nu}$ and $T_{\mu}{}^{\mu}{}T_{\nu}{}^{\nu}$. This pattern is also discernible within a two-dimensional context for the overarching Lagrangian labeled as (3). Consequently, we can deduce the expression for this general perturbation, articulated as the ensuing general flow equation, accurate to the order of $\lambda^{7}$:
@@
$c_{0}=-2a_{2},\,\,\,\,\,c_{1}=-\frac{16{a_{2}}^{3}+3a_{7}{}}{16{a_{2}}^{2}},\,\,\,\,\,$ (3.8)
$c_{2}=-\frac{50176{a_{2}}^{6}+1664{a_{2}}^{3}a_{7}{}-27{a_{7}}^{2}-20a_{2}{}a_{13}{}}{768{a_{2}}^{5}},$
$c_{3}=-\frac{121044992{a_{2}}^{9}+4264960{a_{2}}^{6}a_{7}{}-40512{a_{2}}^{3}{a_{7}}^{2}}{49152{a_{2}}^{8}}$
$-\frac{567{a_{7}}^{3}-38656{a_{2}}^{4}a_{13}{}+720a_{2}{}a_{7}{}a_{13}{}+168{a_{2}}^{2}a_{21}{}}{49152{a_{2}}^{8}}.$
Flow equation (3.7) provides a general perturbative formulation of flow dynamics in $D=2$, capturing a broad class of theoretical models. with the appropriate exponents of $P_{1}$ and $P_{2}$, as they pertain to PDE (1.10). This particular flow equation aligns with the overarching principles of the general Lagrangian, denoted as equation (3).
@@ 4 General root $T\bar{T}$-like flows in two dimensions
The duality-invariant PDE (1.10) can be simplified by expressing $\mathcal{L}$ as a function of $P_{1}$ and
@@
Upon implementing the variable substitution (4.1), the GSMM Lagrangian in (1.7) is reformulated thus:
@@
It can be explicitly verified that Lagrangian (4.3) satisfies PDE (4.2). A similar approach to the previous section can solve self-duality PDE (4.2) with a series of integer powers of the two independent variables, $P_{1}$ and $\mathcal{P}$. In all the theories we consider, there are only two independent invariants $P_{1}$ and $\mathcal{P}$. All higher-order invariants are combinations of these two. For this purpose, we consider the following generic Lagrangian:
@@
By employing a perturbation approach similar to the previous section, we can solve PDE (4.2) using the general solution (4.4), step by step. These conditions are essential for solving Eq. (4.2) up to the $\lambda$ order in the form of $\Big{(}d_{2}{}\to-\tfrac{1}{2}\sqrt{1+4{d_{1}}^{2}}\Big{)}\,\,\,\,\,,\,\,\,\,\,\,\Big{(}d_{3}\to-\frac{d_{4}{}}{4{d_{1}}}\sqrt{1+4{d_{1}}^{2}}\,\,,\,\,\,d_{5}\to-\frac{d_{1}d_{4}}{\sqrt{1+4{d_{1}}^{2}}}\Big{)}$. In this scenario, only one unknown coefficient $d_{n}$ remains at each order of $\lambda$. The perturbed solutions of PDE (4.2), assuming $d_{1}=-\tfrac{1}{2}\sinh(\gamma)$, up to order of $\lambda^{4}$ are as follows:
$\mathcal{L}(\lambda,\gamma)$ $=$ $-\tfrac{1}{2}\bigl{(}P_{1}\cosh(\gamma)+\mathcal{P}\sinh(\gamma)\bigr{)}+\lambda\,d_{4}\frac{\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{2}}{2\coth(\gamma)}$
$+\lambda^{2}\Bigl{(}\tfrac{1}{3}d_{4}^{2}\bigl{(}\mathcal{P}\mathrm{csch}(\gamma)-2P_{1}\mathrm{sech}(\gamma)\bigr{)}\bigl{(}\mathcal{P}\mathrm{csch}(\gamma)+P_{1}\mathrm{sech}(\gamma)\bigr{)}^{2}$
$+\frac{d_{7}\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{3}}{3\bigl{(}\coth(\gamma)\bigr{)}^{2}}\Bigr{)}$
$+\lambda^{3}\biggl{(}\frac{d_{4}d_{7}\bigl{(}\mathcal{P}\mathrm{csch}(\gamma)-3P_{1}\mathrm{sech}(\gamma)\bigr{)}\bigl{(}\mathcal{P}\mathrm{csch}(\gamma)+P_{1}\mathrm{sech}(\gamma)\bigr{)}^{3}}{2\mathrm{csch}(\gamma)}$
$+\frac{d_{11}\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{4}}{4\bigl{(}\coth(\gamma)\bigr{)}^{3}}$
@@
$\times\Bigl{(}-68P_{1}^{2}+9\mathcal{P}^{2}+4(20P_{1}^{2}+3\mathcal{P}^{2})\cosh(2\gamma)+3(-4P_{1}^{2}+\mathcal{P}^{2})\cosh(4\gamma)$
$-72P_{1}\mathcal{P}\bigl{(}\cosh(\gamma)\bigr{)}^{3}\sinh(\gamma)\Bigr{)}$
@@
The most general Lagrangian, denoted $\mathcal{L}(\lambda,\gamma)$ in (4), can be constructed in the presence of two coupling constants $\lambda$ and $\gamma$, up to the order $\lambda^{4}$. In the above Lagrangian, the coefficients of $d_{i}(\gamma)$ depend exclusively on this dependent function of $\gamma$ . The coefficients dependent on $\gamma$ in this Lagrangian can be systematically determined by application of the root flow equation originating from the structure of the Lagrangian coefficients. For this end, we first obtain the momentum energy tensor of this general integrable theory. The energy-momentum tensor corresponding to the general Lagrangian (4) can be determined as follows:
@@
The functions ${\cal H}(P_{1},\mathcal{P},\lambda,\,\gamma)$ and ${\cal U}(P_{1},\mathcal{P},\lambda,\,\gamma)$ are defined in Appendix (A) , in Eqs. (A.4) and (A.5). To derive a root flow equation, we construct two configurations, $T_{\mu\nu}T^{\mu\nu}$ and $T^{\mu}{}_{\mu}T^{\nu}{}_{\nu}$, from the energy-momentum tensor presented in Eq. (4.6). The configurations for these two structures are provided in Appendix (A), in Eqs. (A) and (A.7). We aim to study the root flow equation within the framework of the general Lagrangian (4). We consider a general form of the root flow equation derived from the irrelevant flow equation (3.7). Consequently, the general Lagrangian (4) must satisfy the following flow equation:
@@
The general Lagrangian (4) simplifies to the (SMM) theory in (1.6) in the limit of $\lambda=0$. Notably, there is no $\gamma$-dependent coefficient. The solution to root flow equation (4.7) is derived with the constant $e_{0}=\tfrac{1}{2}$. In the next order of $\lambda$, we encounter the unknown constant $d_{4}$, which determines the $\gamma$-dependence of $d_{4}$ from the solution of root flow equation (4.7). To achieve this, we substitute (A), (A.7), and (4) into root flow equation (4.7) and simplify in the order of $\lambda$, resulting in the following differential equation:
$2\cosh(2\gamma){d}_{4}-\sinh(2\gamma){d}^{\prime}_{4}=0.$ (4.8)
Solving differential equation (4.8) yields the $\gamma$-dependence of the coefficient $d_{4}$ as follows:
@@
where $n_{1}$ is a constant independent of $\gamma$. By applying this method to the order of $\lambda^{2}$, solving root flow equation (4.7) results in a differential equation that depends on the coefficient $d_{7}$ and ${d}^{\prime}_{7}$. The differential equation is as follows:
$n_{1}^{2}\bigl{(}16\cosh(2\gamma)+3(1+4e_{1})\cosh(4\gamma)-51-12e_{1}\bigr{)}$
$+d_{7}\bigl{(}64\mathrm{sech}(\gamma)-96\cosh(\gamma)\bigr{)}+32{d}^{\prime}_{7}\sinh(\gamma)=0.$ (4.10)
By solving differential equation (4), we obtain the $\gamma$-dependence of the coefficient $d_{7}$ as follows:
@@
${d}_{11}$ $=$ $\tfrac{1}{2}n_{1}^{3}\cosh(\gamma)\bigl{(}3(1+4e_{1})\gamma\cosh(3\gamma)$
$-8\sinh(\gamma)\bigr{)}+\bigl{(}n_{3}+n_{1}^{3}(\gamma+4e_{1}\gamma)^{2}\bigr{)}\bigl{(}\cosh(\gamma)\bigr{)}^{3}\sinh(\gamma)$